# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

### v1

@@ 6.1 Elicitation of Human Responses
We split the 10,546 probes from Table 1 into 276 subsets of 38 probes on average, making sure that no subset contained more than one probe with the same relation and the same target word. The time limit for responding to each probe was three minutes. Four participants were assigned to each subset. Participants answered 22 subsets on average.
We asked participants to type up to five relata for each probe. We instructed them that they should use nouns, but no multi-word expressions. We further told participants that the relata could start with either a consonant or a vowel. In addition to the real probes, we used three additional bogus probes (such as “The earth rotates around the [v]) per subset, and rejected subsets where the bogus item was not answered correctly (in this case, only “sun” was accepted). All participants correctly answered all bogus probes, so we were able to accept all responses.
The humans responded with 93,120 word tokens (7,216 word types) in total, of which 10,895 word tokens (18%) are OOR (3,691 OOR types, 51%). 2,106 response lists consist solely of OOR words (20%). On average, we find the first non-OOR words at rank 1.5 in a response list.
On the basis of all responses including OOR, we calculate soundness, completeness, symmetry for antonymy and synonymy (but not for asymmetry), prototypicality, and distinguishability. Our metrics will penalize agents for responding with an OOR word in each case.
@@ 6.2 Response Entropy Analysis
Table 3 lists the ratio of such responses, along with the average number of words in responses. According to this metric, antonymy shows by far the strongest prototypicality at 6.11% of all responses, followed by synonymy at 2.44%. We can observe that meronymy and hyponymy almost never show responses with the strongest prototypicality (only once for meronymy and twice for hyponymy). This might be related to the fact that these two relations happen to also have more relata in the human responses (more than seven words on average) than other relations, which have an average of around five.
Figure 2 shows the distributions of response entropies $\mathcal{R}$ across relations [^15] Distributions are significantly different from each other for all relation pairs except for holonymy-hypernymy and holonymy-synonymy relation pairs, as established by Mann-Whitney U tests with $\alpha$=0.05.. For visibility reasons, we excluded zero scores from Figure 2161616However, zero scores were not excluded when applying the statistical tests..
Figure 2: Response Entropy (excluding zero cases). Boxes enclose second and third quantiles, with the mean shown as orange lines. A lower score shows a stronger prototypicality.
We can see that the strongest prototypicality effect by far is again observed for antonymy, at a mean of 0.83. Synonomy shows the second strongest prototypicality with a mean of 0.90, with hypernymy and holonymy somewhat less prototypical. Hyponymy and meronymy are again at the other extreme, with $\mathcal{R}$ means above 0.95, close to the maximum of 1, and standard deviations lower than those of other relations. This means that most distributions for these two relations are close to uniform. Combining these observations, we conclude that antonymy shows a strong prototypicality effect, but there is hardly any prototypicality effect for hyponymy and meronymy. For our planned prototypicality evaluation of model, we will not use hyponymy and meronymy relations, but only hypernymy, holonymy, antonymy, and synonymy.
If a human response shows a $\mathcal{R}$ of 1, the probe is unable to elicit any prototypical response. We also remove probes whose responses show a $\mathcal{R}$ of 1.0 from these four relations in the evaluation of prototypicality. We find 401 such cases for hypernymy, 199 for hyponymy, 93 for synonymy, 79 for holonymy, 65 for meronymy, and 15 for antonymy. In addition, words that are out of model’s vocabularies will introduce an atrificial deflation in evaluation. We therefore further discard responses that include any word that is not in the intersection vocabulary of the models tested. This results in a total of 4,331 responses which we can use in our prototypicality experiments: 2,406 for hypernymy, 805 for holonymy, 448 for antonymy, and 672 for synonymy.
@@ 7 Model Experiment
For soundness, completeness, and distinguishability, we use all 10,546 probes from Table 1. For model prototypicality, we use 4,331 probes, as described above. For symmetry and asymmetry, we use 17,608 probes. The number is higher than 10,546 because of the trick probes we added (cf. Section 5.2). For symmetry and asymmetry, the values of $k$ we use are 1, 5, and 10.
@@ 7.1 Target Models
We use BERT as one of the MLMs because it is widely used in previous work (we use the cased version). In addition to BERT, we also chose RoBERTa [33] and ALBERT [29], as this allows us to quantify the effect of training objectives and architectures.
In this paper, we also present experiments with CLMs. The most important condition for comparability between neural models is that their vocabularies are as similar as possible. To select the most suitable model, we experimentally determine the Jaccard similarity between two models’ vocabulary. The average Jaccard similarity within the set of MLMs (BERT-RoBERTa, BERT-ALBERT, ALBERT-RoBERTa) is 0.40. We then compare the similarity between each of the two CLMs (OPT and Llama-2) with the three MLMs. When we consider the similarity of OPT with the three MLMs, the average Jaccard similarity (within the new set of four models) increases from 0.40 to 0.50171717Part of the reason for this increase is the fact that OPT and RoBERTa share vocabularies.. For Llama-2, the pairwise similarity drops from 0.40 to 0.32. Based on these results, we choose OPT over Llama-2.
In order to study how model size impacts the learning of semantic relations, we also use models of different sizes within each model family. For BERT and RoBERTa, the two variants we use are of similar sizes, but for OPT, the range of sizes we experiment with is much larger, namely three orders of magnitude. For ALBERT, the size range is somewhere in the middle.
@@
Abbr. Model #Parameters Vocabulary Size Pretraining Set Size
B1 BERT-base 110M 28,996 16GB
B2 BERT-large 340M 28,996 16GB
A1 ALBERT-base 12M 30,000 16GB
A2 ALBERT-large 18M 30,000 16GB
A3 ALBERT-xlarge 60M 30,000 16GB
A4 ALBERT-xxlarge 235M 30,000 16GB
R1 RoBERTa-base 125M 50,265 160GB
R2 RoBERTa-large 355M 50,265 160GB
O1 OPT-125M 125M 50,265 800GB
O2 OPT-350M 350M 50,265 800GB
O3 OPT-1.3B 1.3B 50,265 800GB
O4 OPT-2.7B 2.7B 50,265 800GB
O5 OPT-6.7B 6.7B 50,265 800GB
O6 OPT-13B 13B 50,265 800GB
O7 OPT-30B 30B 50,265 800GB
O8 OPT-66B 66B 50,265 800GB
@@ 7.2 Dealing with Determiner Bias
Some probes require indefinite determiners before the nouns that are to be predicted. Keeping determiners fixed in these probes would introduce a bias towards words with an initial vowel or consonant. For each probe that requires indefinite determiners, we therefore probe the models twice, once with “an” inserted into the probe and the other time with “a” inserted into the probe. From the responses, we create a new distribution over the vocabulary, which is the weighted sum of the two distributions yielded by the two probes. The weights correspond to the relative frequencies of “an” and “a” in the Corpus of Contemporary American English [13]. This synthetic distribution can then be treated as the tested model’s prediction to the probe.
@@ 7.3 Statistical Test
Throughout this study, statistical differences in soundness, symmetry, and asymmetry metrics have been tested using McNemar’s test with $\alpha=0.05$, as these metrics are binary for each target word or each tuple. For completeness and prototypicality, the Wilcoxon signed rank test is used, with $\alpha$=0.05. For distinguishability, no known test exists so we do not test for significance.
@@ 8 Results and Analyses
Before presenting the results of each metric, we examine the general characteristics of responses from each agent. Models produce more OOR responses than humans. The best model, RoBERTa-large (R2), returns the first non-OOR word on average at rank 40.6; recall that this number was 1.5 for humans. Even for the best model, 64% of response lists consist solely of OOR words (all other models’ numbers are even higher). The number for humans was 20%. However, some of the non-OOR responses we received both from models and from humans might be due to holes in our gold standard.
Disregarding OOR responses, Figure 3 shows the distribution of the relation between the first non-OOR word returned by agents and the target word, compared to the gold standard (the prompted-for relation).
Figure 3: The Distribution of the relations that the first non-OOR word is in with the target word for each model family, disregarding all OOR responses.
Humans’ distribution closely resembles the gold standard distribution. However, they slightly underestimate the proportions of holonymy and meronymy. Models’ distributions are less similar to the gold standard, with a clear tendency to underestimate the proportion of hypernymy and meronymy and overestimate the proportion of antonymy. The distribution of relations that is most different from the gold distribution comes from OPT, which performs far worse than the three MLMs. We now present a more detailed analysis as made possible by our specialized metrics.
@@ 8.1 Soundness
Figure 4 shows the result of the soundness evaluation for our models, along with the human ceiling.
Figure 4: Results for soundness.
In the graphs, we show significance using dotted lines. A dotted line indicates that the test established statistically significant differences between all models above and below the line [^18] Please note that this is not equivalent to saying that all models between neighbouring dotted lines are statistically indistinguishable. This may or may not be the case for any pair; the notation we use is a simplification in that it cannot express this aspect..
Humans (H) show high performance on antonymy ($\mathcal{S}=0.90$), with lower scores for the other relations ($0.750.63<\mathcal{S}<$ 0.75, with an average of 0.66). The performance of all models remains far below that of humans: for most relations, even the best model score is less than half the human score.
As for individual models, we can see that for all relations RoBERTa-large (R2) is either the significantly best model, or in the best-performing group. The one relation where models perform relatively well is antonymy, where the best models achieve $\mathcal{S}>0.45$, whereas in other relations best values typically lie around $\mathcal{S}=0.25$. CLMs overall perform less well, with the best scores achieved by any CLM ranging from $\mathcal{S}=0.04$ for synonymy to $\mathcal{S}=0.38$ for antonymy.
@@ 8.2 Completeness
Figure 5 shows the results for completeness.
Figure 5: Results for completeness.
Again, we observe a large gap between models and humans, as was the case for soundness earlier. For relations except for antonymy, the human $\mathcal{C}$ scores range from 0.38 to 0.49 and are therefore overall much lower than the human $\mathcal{S}$ scores, where even the lowest score was above 0.50.
$\mathcal{C}$ scores for models remains below 0.25 for the five relations except antonymy. All other trends are similar to those for soundness: antonymy stands out again as a relation with high completeness, for both humans and models. The overall best performer is again RoBERTa-large (R2).
Among all relations evaluated, antonymy stands out. Both $\mathcal{C}$ and $\mathcal{S}$ scores of antonymy are higher than other relations, for all agents evaluated.
We conclude from the results for OOR-rate, soundness, and completeness that the models only acquire a limited ability to perform relata prediction, which is far below the human ceiling.
8.3 Symmetry and Asymmetry
Let us consider the two symmetric relations first. For $k$=5, Figure 6(a) shows the evaluation results for $\mathcal{M}$ scores [^19] Results for $k$=1 and $k$=10 are given in Appendix B. Overall, we observed similar trends for all values of $k$..
(a) Symmetry.
(b) Asymmetry.
Figure 6: Results for symmetry and asymmetry.
The humans achieve $\mathcal{M}$=0.88 for antonymy and $\mathcal{M}$=0.65 for synonymy. The models similarly perform better for antonymy than they do for synonymy. For antonymy, the best-perfoming group includes BERT-large (B2; $\mathcal{M}=0.62$) and RoBERTa-large (R2; $\mathcal{M}=0.6$)). For synonymy, it includes ALBERT-xxlarge (A4; $\mathcal{M}=0.30$), BERT-large (B2; $\mathcal{M}=0.28$) and RoBERTa-large (R2; $\mathcal{M}=0.27$). Again, for both symmetric relations, the best CLM is left far behind.
We now move to the results of asymmetric relations. Figure 6(b) shows the evaluation results for $\mathcal{A}$ scores, again for $k$=5. The models generally perform better on hypernymy than they do on hyponymy. They also perform better on holonymy than on meronymy. Across all relations, RoBERTa-large (R2) consistently ranks as the top model or is in the best-performing group. However, no results above $\mathcal{A}$=0.40 were measured.
In conclusion, all models tested recognized the asymmetry of four asymmetric relations, and the symmetry of synonymy, only to a limited extent. In contrast, the best set of models recognized the symmetry of antonymy to a relatively high extent, approaching human performance.
@@ 8.4 Prototypicality
Figure 7 gives the evaluation results for $\mathcal{P}$ scores. Remember that prototypicality is not reported for hyponymy and meronymy, as we have established in Section 4 that the human responses showed no prototypicality effects for these relations.
@@
All models for hypernymy, holonymy, and synonymy achieved results below $\mathcal{P}$=0.3. The best results ($\mathcal{P}$=0.26 for holonymy, $\mathcal{P}$=0.23 for synonymy and $\mathcal{P}$=0.21 for hypernymy), were obtained by RoBERTa-large (R2). Antonymy again is the positive outlier relation, with best results ranging around $\mathcal{P}$=0.40 (achieved by ALBERT-xxlarge (A4), RoBERTa-base (R1), BERT-large (B2) and RoBERTa-large (R2)). Therefore, we can conclude that the models only learn a limited degree of prototypicality for hypernymy, holonymy, and synonymy, while they perform relatively well for antonymy.
@@ 8.5 Distinguishability
Figure 8 shows the distinguishability curves for humans and models, with corresponding AuDC values in the lower part of the figure. We can see in the human results that as $p$ increases, the theoretical maximum of $\eta(p;m)$=30 is kept up until after $p=0.5$. The curve then gradually descends, and reaches zero around $p=0.9$. The curves for the models show a faster descent than that for humans; zero $\eta(p;m)$ is reached around $p=0.65$. This means that the upper bound for models only slightly exceeds the lower bound for humans, suggesting a substantial difference between models and humans in their ability to distinguish relation pairs.
@@
Now let us move on to the lower part of Figure 8, which shows the AuDC values. The AuDC tells us how many relation pairs out of 30 are distinguished by an agent, on average. For humans, the $\mathrm{AuDC}$ value is 21.3, whereas models are only able to distinguish 4.0 to 9.8 pairs, fewer than a third of all relations. We can therefore safely say that the meta-relational knowledge of all models we tested is unsatisfactory.
@@
Plotting all pairwise $\mathcal{D}$ scores for an agent, we can create the agent’s distinguishability matrix. This serves to examine more deeply which relations are mistaken for which other ones. The relation given at the row position is the prompted relation, and the relation given at the column position is the relation the model responded with. The depth of grey shade expresses the degree of distinguishability, lighter cells indicating that the row relation is more often confused with the column relation.
@@
Figure 9 presents $\mathcal{D}$ scores for entire model families, calculated as averages over all model variations of MLM (left) and CLM (middle), in comparison to the human ceiling (right) [^20] The complete list of distinguishability matrices for all model variants is provided in Appendix C.. Humans, despite their generally good ability to distinguish semantic relations, tend to confuse hyponymy and hypernymy with synonymy, as can be seen from the lighter cells at the leftmost bottom ((SYN, HYP) and (SYN, HPO)) and at the rightmost top ((HYP, SYN) and (HPO, SYN)). This agrees with findings by Chaffin and Clark: humans perceive synonymy as being close to hypernymy and hyponymy (cf. Section 2). We also observe that humans’ $\mathcal{D}$ scores for hypernymy versus both holonymy and meronymy (lighter cells, (HOL, HYP) and (MER, HYP) in the leftmost middle) are relatively low as well. This aligns with the theoretical assumption by Cruse [12], Winston et al [64] and Joosten [27] that there are similarities between holonymy and hypernymy.
@@
What stands out in both the models’ and humans’ distinguishability matrices are the high scores when antonymy is the relation prompted (visible in the antonymy row). However, in the antonymy column, which shows the cases where the model responded with antonyms although the prompt was associated with a non-antonymy relation, high $\mathcal{D}$ scores are observed only for humans, but not for models.
To sum up, we observe that models can distinguish antonymy from other relations, but not other relations from antonymy. Antonymy is also the relation for which the models show the highest scores in all metrics we introduced: not only for distinguishability, but also for soundness, completeness, symmetry, and prototypicality. These observations lead us to conclude that there must be some kind of antonymy bias: irrespective of the relation prompted, models consistently prefer antonyms of the target words. We now show an example of this antonymy bias.
@@
Table 5: Top three words in the prediction of models given the probe “an answer is similar to [det] [v]. Antonyms are shown in boldface.
Abbr. Agent Top 1 Top 2 Top 3
H Humans response reply solution
B1 BERT-base question answer statement
B2 BERT-large question statement answer
R1 RoBERTa-base question yes answer
R2 RoBERTa-large question answer argument
A1 ALBERT-base question query error
A2 ALBERT-large question answer query
A3 ALBERT-xlarge question inquiry query
A4 ALBERT-xxlarge question reply query
[…]
@@ H Humans response reply solution
Humans correctly produced only synonyms for this probe, but all models except for the 4 smallest OPT models return “question” as the first response; O1, O3 and O4 return “question” as their first non-OOR response, while OPT-350M (O2) returns only non-OOR words in the top three. This suggests that the relation of “question” and “answer” is learned firmly by the models, particularly the MLMs.

### v4

@@ 6.1 Elicitation of Human Responses
We split the 10,507 probes from Table 1 into 276 subsets of 38 probes on average, making sure that no subset contained more than one probe with the same relation and the same target word. The time limit for responding to each probe was three minutes. Four participants were assigned to each subset. Participants answered 22 subsets on average.
We asked participants to type up to five relata for each probe. We instructed them that they should use nouns, but no multi-word expressions. We further told participants that the relata could start with either a consonant or a vowel. In addition to the real probes, we used three additional bogus probes (such as “The earth rotates around the [v]”) per subset, and rejected subsets where the bogus item was not answered correctly (in this case, only “sun” was accepted). All participants correctly answered all bogus probes, so we were able to accept all responses.
The humans responded with 88,896 word tokens (7,077 word types) in total, of which 10,699 word tokens (12.0%) are OOR (3,599 OOR types, 51.0%). 396 response lists consist solely of OOR words (3.8%). On average, we find the first non-OOR words at rank 1.5 in a response list.
On the basis of all responses including OOR, we calculate soundness, completeness, symmetry for antonymy and synonymy, prototypicality, and distinguishability. Our metrics will penalize agents for responding with an OOR word in each case.
@@ 6.2 Response Entropy Analysis
Table 3 lists the ratio of such responses, along with the average number of words in responses. According to this metric, antonymy shows by far the strongest prototypicality at 6.11% of all responses, followed by synonymy at 2.44%. We can observe that meronymy and hyponymy rarely show responses with the strongest prototypicality (only once for meronymy and twice for hyponymy). This might be related to the fact that these two relations happen to also have more relata in the human responses (more than seven words on average) than other relations, which have an average of around five.
Figure 2 shows the distributions of response entropies $\mathcal{R}$ across relations [^13] Distributions are significantly different from each other for all relation pairs except for holonymy-hypernymy and holonymy-synonymy relation pairs, as established by Mann-Whitney U tests with $\alpha$=0.05..
Figure 2: Response Entropy. A lower score shows a stronger prototypicality. Boxes enclose second and third quantiles, with the mean shown as orange lines.
We can see that the strongest prototypicality effect by far is again observed for antonymy, at a mean of 0.83. Synonymy shows the second strongest prototypicality with a mean of 0.90, with hypernymy and holonymy somewhat less prototypical. Hyponymy and meronymy are again at the other extreme, with $\mathcal{R}$ means above 0.95, close to the maximum of 1, and standard deviations lower than those of other relations. This means that most distributions for these two relations are close to uniform. Combining these observations, we conclude that antonymy shows a strong prototypicality effect, but there is hardly any prototypicality effect for hyponymy and meronymy. For our planned prototypicality evaluation of the model, we will not use hyponymy and meronymy relations, but only hypernymy, holonymy, antonymy, and synonymy.
If a human response shows a $\mathcal{R}$ of 1, the probe is unable to elicit any prototypical response. We also remove probes whose responses show a $\mathcal{R}$ of 1.0 from these four relations in the evaluation of prototypicality. We find 399 such cases for hypernymy, 197 for hyponymy, 93 for synonymy, 79 for holonymy, 65 for meronymy, and 15 for antonymy. In addition, words that are out of the model’s vocabulary will introduce an artificial deflation in evaluation. We therefore further discard responses that include any word that is not in the intersection vocabulary of the models tested. This results in a total of 4,282 responses which we can use in our prototypicality experiments: 2,375 for hypernymy, 798 for holonymy, 447 for antonymy, and 662 for synonymy.
@@ 7 Model Experiment
For soundness, completeness, and distinguishability, we use all 10,507 probes from Table 1. For model prototypicality, we use 4,282 probes, as described above. The number is higher than 10,507 because of the trick probes we added (cf. Section 5.2). For symmetry, we use 1, 5, and 10 for the value of $k$.
@@ 7.1 Target Models
[…]
In order to study how model size impacts the learning of semantic relations, we also use models of two different sizes within each model family. For BERT and RoBERTa, the two variants we use are of similar sizes, but for Llama, the range of sizes we experiment with is much larger, namely one order of magnitude.
@@
Abbr. Model #Parameters
Vocabulary
Size
Pretraining Corpus Size
(in token numbers)
B1 BERT-base 110M 28,996 3.3B
B2 BERT-large 340M 28,996 3.3B
R1 RoBERTa-base 125M 50,265 30B
R2 RoBERTa-large 355M 50,265 30B
L1 Llama-8B 8B 128,256 70T
L2 Llama-70B 70B 128,256 70T
@@ 7.2 Dealing with Determiner Bias
Some probes require indefinite determiners before the nouns that are to be predicted. Keeping determiners fixed in these probes would introduce a bias towards words with an initial vowel or consonant. For each probe that requires indefinite determiners, we, therefore, probe the models twice, once with “an” inserted into the probe and the other time with “a” inserted into the probe. From the responses, we create a new distribution over the vocabulary, which is the weighted sum of the two distributions yielded by the two probes. The weights correspond to the relative frequencies of “an” and “a” in the Corpus of Contemporary American English [16]. This synthetic distribution can then be treated as the tested model’s prediction of the probe.
@@ 7.3 Statistical Test
Throughout this study, statistical differences in soundness, and symmetry metrics have been tested using McNemar’s test with $\alpha=0.05$, as these metrics are binary for each target word or each tuple. For completeness and prototypicality, the Wilcoxon signed rank test is used, with $\alpha$=0.05. For distinguishability, no known test exists so we do not test for significance.
@@ 8 Results and Analyses
We present results and analyses based on them. Before presenting the results of each metric, we examine the general characteristics of responses from each agent by evaluating OOR words.
8.1 OOR Words in Responses
[…]
@@ 8.2 Soundness
Figure 3 shows the result of the soundness evaluation for the models, along with the human ceiling.
Figure 3: Results of soundness.
In the graphs, we show significance using dotted lines. A dotted line indicates that the test established statistically significant differences between all models above and below the line [^15] Note that this is not equivalent to saying that all models between neighbouring dotted lines are statistically indistinguishable. This may or may not be the case for any pair; the notation we use is a simplification in that it cannot express this aspect..
Humans (H) show high performance on antonymy ($\mathcal{S}=0.90$), with lower performance for the other relations ($0.54<\mathcal{S}<0.75$, with an average of 0.66). The performance of all models remains far below that of humans: for most relations, even the best model score is less than half the human score.
All models perform relatively well for antonymy, where the best model Llama-8B (L1) achieves $\mathcal{S}=0.57$, whereas for other relations, best values typically lie around $\mathcal{S}=0.30$.
When comparing CLMs and MLMs, we can see that CLMs (L1 and L2) outperform MLMs (B1, B2, R1, and R2) for most relations. For hypernymy, holonymy, antonymy, and synonymy, either Llama-8B or Llama-70B outperforms all the MLMs. MLMs have an advantage in hyponymy and meronymy. RoBERTa-large (R2) leads all other models with a score of around $\mathcal{S}=0.25$, which is significantly better than that of all other models.
@@ 8.3 Completeness
Figure 4 shows the results of completeness.
Figure 4: Results of completeness.
Again, we observe a large gap between models and humans, as was the case for soundness. For relations except for antonymy, the human $\mathcal{C}$ scores range from 0.38 to 0.50 and are therefore overall much lower than the human $\mathcal{S}$ scores, where even the lowest score was above 0.50.
The models’ $\mathcal{C}$ scores remain below 0.25 for the five relations except for antonymy. All other trends are similar to those for soundness: antonymy stands out again as a relation with high completeness, for both humans and models. CLMs’ priority can be observed for hypernymy, holonymy, and antonymy but not for hyponymy and meronymy. For these two relations, the normally observed priority of CLMs is often even reversed: all MLMs outperform CLMs for hyponymy and three out of four MLMs do so for meronymy. For synonymy, RoBERTa-large (R2) shows significant advantage than Llama-8B (L1) in $\mathcal{C}$ scores, whereas their difference is not significant in $\mathcal{S}$ scores.
We conclude from the results for the OOR rate, soundness, and completeness that the models only acquire a limited ability to perform relata prediction, which is far below the human ceiling.
8.4 Symmetry
[…]
Figure 5: Symmetry.
The humans achieve $\mathcal{M}$=0.88 for antonymy and $\mathcal{M}$=0.65 for synonymy. The models similarly perform better for antonymy than they do for synonymy. For both relations, the best-perfoming model remains Llama-70B (L2; $\mathcal{M}=0.67$ for antonymy and $\mathcal{M}=0.34$ for synonymy). While CLMs generally keep the leading position for antonymy symmetry, MLMs are not generally inferior to CLMs for synonymy symmetry. BERT-large (B2; $\mathcal{M}=0.28$) and RoBERTa-large (R2; $\mathcal{M}=0.27$) achieve significantly better results than Llama-8B (L1, $\mathcal{M}=0.21$).
In conclusion, all models tested recognize the symmetry of synonymy only to a limited extent. In contrast, the best model recognizes the symmetry of antonymy to a relatively high extent, approaching human performance.
@@ 8.5 Prototypicality
Figure 6 gives the evaluation results for the $\mathcal{P}$ scores. Remember that prototypicality is not reported for hyponymy and meronymy, as we have established in Section 4 that the human responses showed no prototypicality effects for these relations.
@@
For hypernymy, holonymy, and synonymy, all models achieve a $\mathcal{P}$ score below 0.3. The best-performing models for each relation are as follows. For synonymy, they are Llama-70B (L2; $\mathcal{P}=0.26$) and RoBERTa-large (R2; $\mathcal{P}=0.22$); for holonymy, RoBERTa-large (R2; $\mathcal{P}=0.25$), two Llama variants, and BERT-large (L1, L2, B2; all $\mathcal{P}=0.23$); for hypernymy, Llama-8B (L1; $\mathcal{P}=0.22$).
Antonymy again is the positive outlier relation, with best results ranging around $\mathcal{P}=0.40$, a value achieved by all models except for BERT-base (B1).
In summary, these results indicate that models capture only a limited degree of prototypicality for hypernymy, holonymy, and synonymy, but perform relatively well on antonymy.
@@ 8.6 Distinguishability
Figure 7 shows the distinguishability curves for humans and models, with corresponding AuDC values in the lower part of the figure. We can see in the human results that as $p$ increases, the theoretical maximum of $\eta(p;m)$=30 is kept up until after $p=0.5$. The curve then gradually descends and reaches zero around $p=0.9$. The curves for the models show a faster descent than that for humans; zero $\eta(p;m)$ is reached around $p=0.65$ for MLMs and around $p=0.75$ for CLMs. This means that, although CLMs outperform MLMs, the upper bound for all models remains far below that for humans, suggesting a substantial difference between models and humans in their ability to distinguish relations.
@@
Now let us move on to the lower part of Figure 7, which shows the AuDC values. The AuDC tells us how many relation pairs out of 30 are distinguished by an agent, on average. For humans, the $\mathrm{AuDC}$ value is 21.3, whereas models are only able to distinguish 7.5 (B1) to 12.3 (L2) pairs. It means that models, at best, would distinguish roughly a third of all relation pairs. We can therefore safely say that the distinguishability of all models we tested is unsatisfactory.
@@
Plotting all pairwise $\mathcal{D}$ scores for an agent, we can create the agent’s distinguishability matrix. This serves to examine more deeply which relations are mistaken for which other ones. Figure 8 presents distinguishability matrices for entire model families, calculated as averages over all model variations of MLM (left) and CLM (middle), in comparison to the human ceiling (right) [^17] The complete list of distinguishability matrices for all model variants is provided in Appendix C.. The relation given at the row position is the prompted relation, and the relation given at the column position is the relation the model responded with. The depth of grey shade expresses the degree of distinguishability, with lighter cells indicating that the row relation is more often confused with the column relation.
@@
Humans, despite their generally good ability to distinguish semantic relations, tend to confuse hyponymy and hypernymy with synonymy, as can be seen from the lighter cells at the leftmost bottom ((SYN, HYP) and (SYN, HPO)) and at the rightmost top ((HYP, SYN) and (HPO, SYN)). This agrees with findings by Chaffin and Clark: humans perceive synonymy as being close to hypernymy and hyponymy (cf. Section 2). We also observe that humans’ $\mathcal{D}$ scores for hypernymy versus both holonymy and meronymy (lighter cells, (HYP, HOL) and (HYP, MER) in the leftmost middle) are relatively low as well. This aligns with the theoretical assumption by Cruse [15], Winston et al [69] and Joosten [31] that there are similarities between holonymy and hypernymy.
@@
What stands out in both the models’ and humans’ distinguishability matrices are the high scores for antonymy (visible in the antonymy row). When the prompt is associated with antonymy, humans and models are both able to rank antonyms of the target word before other $r$-relata.
Let us now look at the antonymy column in the MLM’s matrix. The antonymy column shows the extent to which the model responds with antonyms although the prompt was associated with a non-antonymy relation. We can now see a contrasting pattern: for each row, the $\mathcal{D}$ scores in the antonymy column are the lowest of all relations. Given these results, we can conclude from this that MLMs generally prefer antonyms of target words, irrespective of the relation prompted.
In contrast, CLMs generally show a weaker preference toward antonyms than MLMs. For CLMs, it is never the case that the lowest $\mathcal{D}$ scores appear in the antonymy column. However, for synonymy prompts, CLMs show mild confusion with antonymy ($\mathcal{D}=0.32$, the second lowest score in synonymy row). CLMs tend to struggle with a different distinction, namely that between hyponymy and hypernymy on the one hand, and between hyponymy and synonymy on the other. This can be seen from the fact that the respective $\mathcal{D}$ scores are below 0.20 (lighter cells, (HYP, HPO) and (SYN, HPO) in the second row).
In summary, the only distinction that the models were able to perform well is that between antonymy and other relations. Concerning the distinctions between non-antonymy relations, all models perform poorly. If confronted with an antonymy probe, the models are able to correctly produce antonyms. However, if confronted with a non-antonymy probe, in some cases, the models even make the mistake of incorrectly producing antonyms. We now show an example of this confusion.
@@
Table 5: Top three words in the prediction of models given the probe “an answer is similar to [det] [v]”
. Antonyms are shown in boldface. Abbr. Agent Top 1 Top 2 Top 3 H Humans response reply solution B1 BERT-base question answer statement B2 BERT-large question statement answer R1 RoBERTa-base question yes answer R2 RoBERTa-large question answer argument L1 Llama-8B question in solution L2 Llama-70B [ but question
Table 5 shows the different agents’ top three relata in response to the probe “an answer is similar to [det] [v]”. The probe is expected to elicit synonyms of “answer”: “response”, “result”, “solution”, “reply” and “resolution”, according to our relatum set. Antonyms of “answer”, such as “question”, are shown in boldface.
@@
Humans correctly produced only synonyms for this probe, but all models return “question” as the first non-OOR response. All models except for Llama-70B (L2) return “question” as their first response. This suggests that the relation of “question” and “answer” is learned firmly by the models, particularly the MLMs, leading to the problems of confusion between synonymy and antonymy.
Summing up all the observations so far, we can conclude that antonymy is the relation for which the models show the highest scores in all metrics we introduced: not only for distinguishability, but also for soundness, completeness, symmetry, and prototypicality. These lead us to conclude that there must be antonymy bias: models learn only antonymy to a relatively good extent.